If the product of -3 and the opposite of a number is decreased by 7, the result is 1 greater than the number. What is the number?
Answer by ankor@dixie-net.com(22740)

If the product of -3 and the opposite of a number is decreased by 7, the result is 1 greater than the number. What is the number?
:
let n = the number
then
-n = the opposite of that number
:
The equation from the statement given:
(-3*-n) - 7 = n + 1
3n - 7 = n + 1
3n -n = 1 + 7
2n = 8
n = 8/2
n = 4 is the number